OpenStudy (anonymous):

Prove without using calculator, [5^(1/2)+2]^4+[5^(1/2)-2]^4=322 someone pls help me....

OpenStudy (amistre64):

at all amounts to reading the notation correctly and applying it

OpenStudy (amistre64):

i dont see any "easy" way around it other than its given to you to prcatice seeing how the math is done

OpenStudy (amistre64):

expand it all out and add it all together; what part are you stuck with?

OpenStudy (anonymous):

without using calculator...

OpenStudy (amistre64):

we could prolly clean it up a bit using abstract notations; but in the end the math still has to be done

OpenStudy (anonymous):

use \[(a+b)^4=a^4+4a^3b+4a^2b^2+4ab^3+b^4\]??

OpenStudy (anonymous):

that should be a\[6a^2b^2\]

OpenStudy (amistre64):

(A+n)^4 + (A-n)^4 = 322 (A^2+ 2An +n^2)^2 + (A^2 - 2An +n^2)^2 A^2+ 2An +n^2 A^2+ 2An +n^2 -------------- A^4 +A^3(2n) + (An)^2 +A^3(2n) +4(An)^2 +2A n^3 + (An)^2 +2A n^3 + n^4 ------------------------------------ A^4 + 2nA^3 + 6(An)^2 + 4A n^3 + n^4 yeah, binomial expansion might be quicker

OpenStudy (anonymous):

ok... maybe i calculate wrongly...

OpenStudy (anonymous):

One thing about conjugates if that if \[(a+b\sqrt{c})^n = x+y\sqrt{c}\] then: \[(a-b\sqrt{c})^n = x-y\sqrt{c}\]

OpenStudy (amistre64):

and since the next one has -n we simply switch odds powers A^4 + 4A^3 n + 6(An)^2 + 4A n^3 + n^4 A^4 - 4A^3 n + 6(An)^2 - 4A n^3 + n^4 ------------------------------------ 2A^4 +12(An)^2 +2n^4

OpenStudy (anonymous):

ok i get it... is \[(a+b)^4=a^4+4a^3b+4a^2b^2+4ab^3+b^4\]

OpenStudy (amistre64):

2(sqrt(5))^4 + 12(2sqrt(5))^2 +2(2)^4 = 322 ??

OpenStudy (amistre64):

2(5.5) + 12(4.5) + 2(4.4) 50 + 240 + 32 = 322 ?? 1 240 50 32 ---- 322